# The modified diagonal cycles of Hypergeometric curves

### v1

@@ Abstract.
For each $N\geq 2$, Asakura and Otsubo have recently introduced a smooth family of algebraic curves $\{X_{N,\lambda}\}_{\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}}$ in characteristic 0 that is closely related to hypergeometric functions and the Fermat curve of degree $N$. In this paper, we study the Gross-Kudla-Schoen modified diagonal 1-cycles of these curves. We prove that if $p\geq 3$ is a prime, then for every $\lambda$ the Griffiths Abel-Jacobi image of the modified diagonal cycle of $X_{p,\lambda}$ is nontrivial for every cuspidal choice of a base point. On the other hand, we show that the modified diagonal cycle and hence the Ceresa cycle of $X_{3,\lambda}$ is torsion in the Chow group for every $\lambda$ and every choice of a base point.
@@
Ceresa cycle; Abel-Jacobi map; Hypergeometric curve.
@@ 1. Introduction
Let $X$ be a smooth projective curve over $\mathbb{C}$ with Jacobian $\operatorname{Jac}(X)$. Denoting $\Delta_{123}=\{(x,x,x)\mid x\in X\}\in\operatorname{CH}_{1}(X^{3})$ where $\operatorname{CH}$ is the Chow group [^1] Unless otherwise indicated, all the Chow groups in the paper are with coefficients in $\mathbb{Z}$., for any point $e$ in $X$ the modified diagonal cycle $\Delta_{{\rm GKS},e}(X)$ of Gross-Kudla-Schoen (first introduced by Gross and Scheon in [11]) is the homologically trivial algebraic 1-cycle [^2] All constructions of the paper involving algebraic cycles factor through the Chow groups. Thus with abuse of terminology, we use the same notation and language for an algebraic cycle and its class in the Chow group.
$\Delta_{{\rm GKS},e}(X):=\Delta_{123}-\Delta_{12}-\Delta_{23}-\Delta_{13}+\Delta_{1}+\Delta_{2}+\Delta_{3},$
where
$\Delta_{12}=\{(x,x,e)\mid x\in X\},\$$\Delta_{23}=\{(e,x,x)\mid x\in X\},\$$\Delta_{13}=\{(x,e,x)\mid x\in X\},$
$\Delta_{1}=\{(x,e,e)\mid x\in X\},\$$\Delta_{2}=\{(e,x,e)\mid x\in X\},\$$\Delta_{3}=\{(e,e,x)\mid x\in X\}.$
Another homologically trivial 1-cycle naturally associated with the pointed curve $(X,e)$ is the Ceresa cycle
@@
where $X_{e}$ is the image of $X$ under the map $X\rightarrow\operatorname{Jac}(X)$ defined by $x]\mapsto[x]-[e]$. Thanks to the works of Colombo and van Geeman [7] and S. W. Zhang [26], the two cycles $\Delta_{{\rm GKS},e}(X)$ and $\operatorname{Cer}_{e}(X)$ are intimately related modulo torsion. In particular, one is torsion in the Chow group if and only if the other is. In addition, one knows the following two general facts about them: First, if $X$ is hyperelliptic and $e$ is a Weierstrass point, then they are torsion [11]. Second, if $X$ is a generic complex curve of genus $\geq 3$, then the two cycles are of infinite order for any base point $e$ [5].
@@
    Infinitude of order in the Chow group: Using the work of Darmon-Rotger-Sols [8] on Chow-Heegner points, the first author and Murty showed in [9] that for every positive integer $N$ that has a prime divisor greater than 7, the Ceresa cycle of the Fermat curve $F(N)$ of degree $N$ is non-torsion in the Chow group for every choice of base point $e$. This work was generalized by the second author in [19] to an infinite number of non-hyperelliptic simple Fermat quotients. Using the same method, Kerr-Li-Qiu-Yang [14] and Lupoian-Rawson [18] have proved that the Ceresa cycles of an infinite number of modular curves are non-torsion in the Chow group.
@@
We prove two main results:
@@
Let $p\geq 3$ be a prime and $e\in X_{p,\lambda}$ be a cusp (i.e. a point satisfying $x_{0}x_{1}y_{0}y_{1}=0$). Then for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$ and positive integer $l\leq(p-1)/2$, we have
@@
where $\Phi_{1}(\Delta_{{\rm GKS},e}(X_{p,\lambda}))$ is the complex Abel-Jacobi image of $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ in the intermediate Jacobian of the integral Hodge structure $H_{3}(X^{3})$. In particular, the modified diagonal cycle $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ is nontrivial in the integral Chow group $\operatorname{CH}_{1}(X^{3})$ for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$.
For a cuspidal choice of a base point $e$ as above, one has $(2(p-1)^{2}-2)[e]=K(X_{p,\lambda})$ in $\operatorname{CH}_{0}(X_{p,\lambda})\otimes\mathbb{Q}$, where $K(X_{p,\lambda})$ is the canonical divisor (see Corollary 3.4). Thus Theorem 1.1 does not follow from the general fact observed in [16, Proposition 2.2] that if the modified diagonal cycle (or the Ceresa cycle) is torsion for a pointed curve $(X,e)$ of genus $g$, then (2g2)[e]=K(X)(2g-2)[e]=K(X) in the Chow group over $\mathbb{Q}$. To our knowledge, Theorem 1.1 provides the first explicit example of a smooth family of curves of arbitrarily large genus where the modified diagonal cycle is nontrivial in the entire family. Since the hypergeometric curve $X_{p,\lambda}$ is defined over $\mathbb{Q}(\lambda)$, in particular, the result gives infinitely many curves of a given genus $(p-1)^{2}$ (for every prime $p$) over any subfield of $\mathbb{C}$ with nontrivial modified diagonal cycles.
Unfortunately, we do not know whether this cycle is non-torsion or not for an arbitrary $p$. In particular, since the known relationships between the Ceresa cycle and the modified diagonal cycle (see [7, §2.9], [26, Theorem 1.5.5]) are modulo torsion, our Theorem 1.1 does not have any implication about the Ceresa cycle. However, for $p=3$ we can prove the following:
@@
For every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$ and every choice of a base point $e\in X_{3,\lambda}$, the modified diagonal cycle $\Delta_{{\rm GKS},e}(X_{3,\lambda})$ (and hence the Ceresa cycle $\operatorname{Cer}_{e}(X_{3,\lambda})$) is torsion in the Chow group.
The curve $X_{3,\lambda}$ is not hyperelliptic (see Proposition 3.1), so that this gives an explicit example of a family of non-hyperelliptic curves of genus 4 with torsion modified diagonal cycles. Another example of such a family has been given by Qiu-Zhang [24]. The proof of Theorem 1.2 uses the cohomological vanishing criteria of Qiu-Zhang [24, Theorem 1.2.1] and Laga-Shnidman [17, Theorem A]. [^4] We use the latter criterion to make the argument shorter, but the former criterion would also suffice in this case.
@@ 2.2. A criterion for finiteness of the order of the modified diagonal cycle
Let $X$ be a smooth projective curve of genus $g\geq 3$ over $\mathbb{C}$. Let $\operatorname{Jac}(X)$ be the Jacobian of $X$. Given a point $e\in X$, let
@@
be the modified diagonal cycle defined in [11] (see the §1 to recall the definition). We refer to this as the Gross-Kudla-Schoen modified diagonal cycle with base point $e$. Gross and Schoen showed that this cycle is homologically trivial [11, Proposition 3.1]. One also has the Ceresa cycle
@@
(see §1). Qui and W. Zhang [24, Theorem 1.21.2] have given a sufficient cohomological criterion for the torshion-ness of $\operatorname{Cer}_{e}(X)$ in $\operatorname{CH}_{1}(\operatorname{Jac}(X))_{\rm hom}$. This criterion was then strengthened by Laga and Shnidman in [17]. By [26, Theorem 1.5.5], the torsion-ness of $\operatorname{Cer}_{e}(X)$ and $\Delta_{{\rm GKS},e}(X)$ in the corresponding Chow groups are equivalent. Thus the criteria of Qui-Zhang and Laga-Shnidman also apply to the modified diagonal cycle. We recall the latter criterion (stated for $\Delta_{{\rm GKS},e}(X)$):
Theorem 2.1 ([17, Theorem A]).
Let $\operatorname{Aut}(X)$ be the automorphism group of $X$. Consider the action of $\operatorname{Aut}(X)$ on $\wedge^{3}H^{1}_{\rm dR}(X)$ through its natural action on each factor of the de Rham cohomology $H^{1}_{\rm dR}(X)$, and let $(\wedge^{3}H^{1}_{\rm dR}(X))^{\operatorname{Aut}(X)}$ be the fixed part under this action. If $(\wedge^{3}H^{1}_{\rm dR}(X))^{\operatorname{Aut}(X)}=0$, then $\Delta_{{\rm GKS},e}(X)$ is torsion.
Note that [17] works with Chow groups with coefficients in $\mathbb{Q}$, hence the result is stated as a vanishing condition in that paper.
@@ 2.3. Hypergeometric curves
The differential form $\omega^{a,b}_{N,\lambda}$ is holomorphic on $X_{N,\lambda}$, and $\eta^{a,b}_{N,\lambda}$ is non-holomorphic but of the second kind. The collection of differential forms $\omega^{a,b}_{N,\lambda}$ and $\eta^{a,b}_{N,\lambda}$ forms a basis of $H^{1}_{\rm dR}(X_{N,\lambda})^{\chi_{N}^{a,b}}$.
@@ 3.1. Determination of the hyperelliptic cases
In this subsection we will determine when a hypergeometric curve or quotient is hyperelliptic. This is important to put Theorems 1.1 and Theorem 1.2 in a better context, since the behaviour of the modified diagonal cycle for hyperelliptic curves is well understood.
@@
Suppose that $X_{N,\lambda}$ is hyperelliptic. Let $\tau$ be its hyperelliptic involution. Since the genus of $X_{N,\lambda}$ is greater than one, $\tau$ lies in the center of the automorphism group of $X_{N,\lambda}$, so that there is a canonical isomorphism
$\operatorname{Aut}(X_{N,\lambda})/\langle\tau\rangle=\operatorname{Aut}(X_{N,\lambda}/\tau)=\operatorname{Aut}(\mathbb{P}^{1}).$
Now first assume that $\tau\not\in\mu_{N}\times\mu_{N}$. Then we have
$\mu_{N}\times\mu_{N}\subset\operatorname{Aut}(X_{N,\lambda})/\langle\tau\rangle=\operatorname{Aut}(\mathbb{P}^{1})\subset\operatorname{PGL}_{2}(\mathbb{C}).$
By the well known characterization of the finite subgroups of $\operatorname{PGL}_{2}(\mathbb{C})$ due to Klein, $\operatorname{Aut}(X_{N,\lambda})/\langle\tau\rangle$ is isomorphic to one of the following: the cyclic group $C_{n}$ of order $n\geq 1$, the dihedral group $D_{n}$ of order $2n$ for $n\geq 2$, the alternating group $A_{4}$ or $A_{5}$, or the symmetric group $S_{4}$. However, these groups do not include any copy of the group $\mu_{N}\times\mu_{N}$.
Secondly, we suppose that $\tau\in\mu_{N}\times\mu_{N}$. Then $N$ is even and $\tau$ is given by one of the three elements of order 2 in $\mu_{N}\times\mu_{N}\subset\operatorname{Aut}(X_{N,\lambda})$, i.e., one of the following:
@@
Next, we consider the quotients of $X_{N,\lambda}$.
@@
[…]

### v2

@@ Abstract.
For each $N\geq 2$, Asakura and Otsubo have recently introduced a smooth family of algebraic curves $\{X_{N,\lambda}\}_{\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}}$ in characteristic 0 that is closely related to hypergeometric functions and the Fermat curve of degree $N$. In this paper, we study the Gross-Kudla-Schoen modified diagonal 1-cycles of these curves. We prove that if $p\geq 3$ is a prime, then for every $\lambda$ the Griffiths Abel-Jacobi image of the modified diagonal cycle of $X_{p,\lambda}$ is nontrivial for every cuspidal choice of a base point. On the other hand, we show that the modified diagonal cycle of $X_{3,\lambda}$ is torsion in the Chow group for every $\lambda$ and cuspidal base divisor.
@@
Ceresa cycle; modified diagonal cycle; Abel-Jacobi map; Hypergeometric curve.
@@ 1. Introduction
Let $X$ be a smooth projective curve over $\mathbb{C}$ with Jacobian $\operatorname{Jac}(X)$. Let $e$ be a divisor of degree 1 on $X$. Denoting $\Delta=\{(x,x)\mid x\in X\}\in\operatorname{CH}_{1}(X^{2})$ where $\operatorname{CH}$ is the Chow group [^1] Unless otherwise indicated, all the Chow groups in the paper are with coefficients in $\mathbb{Z}$. All constructions of the paper involving algebraic cycles factor through Chow groups. Thus with abuse of terminology, we use the same notation and language for an algebraic cycle and its class in the Chow group., the modified diagonal cycle $\Delta_{{\rm GKS},e}(X)$ of Gross-Kudla-Schoen based at $e$ (first introduced by Gross and Schoen in [11]) is the homologically trivial algebraic 1-cycle
$\Delta_{{\rm GKS},e}(X):=\Delta_{123}-\Delta_{12}-\Delta_{23}-\Delta_{13}+\Delta_{1}+\Delta_{2}+\Delta_{3}\in\operatorname{CH}_{1}(X^{3}).$
Here,
$\Delta_{123}=\{(x,x,x)\mid x\in X\},$
$\Delta_{12}=\operatorname{pr}^{*}_{1,2}(\Delta)\cdot\operatorname{pr}_{3}^{*}(e),\$$\Delta_{23}=\operatorname{pr}^{*}_{2,3}(\Delta)\cdot\operatorname{pr}_{1}^{*}(e),\$$\Delta_{13}=\operatorname{pr}^{*}_{1,3}(\Delta)\cdot\operatorname{pr}_{2}^{*}(e),$
$\Delta_{1}=\operatorname{pr}_{2,3}^{*}(e\times e),$ $\Delta_{2}=\operatorname{pr}_{1,3}^{*}(e\times e),\ \quad\quad\$$\Delta_{3}=\operatorname{pr}_{1,2}^{*}(e\times e),$
where $\operatorname{pr}_{i}$ (resp. $\operatorname{pr}_{i,j}$) denotes the projection from $X^{3}$ onto the $i$-th factor (resp. the product of the $i$-th and the $j$-th factors) and $\cdot$ is the intersection product. Another homologically trivial 1-cycle naturally associated with the pair $(X,e)$ is the Ceresa cycle (defined in [5])
@@
where $X_{e}$ is the image of $X$ under the map $X\rightarrow\operatorname{Jac}(X)$ defined by $x\mapsto[x]-e$. Thanks to the works of Colombo and van Geemen [7], S-W. Zhang [27], Laga and Shnidman [16] and very recently Lagarde et al [18], the two cycles $\Delta_{{\rm GKS},e}(X)$ and $\operatorname{Cer}_{e}(X)$ are intimately related. In particular, one is torsion in the Chow group if and only if the other is. In addition, one knows the following two general facts about them: First, if $X$ is hyperelliptic and $e$ is supported on Weierstrass points, then the two cycles are torsion [11]. Second, if $X$ is a generic complex curve of genus $\geq 3$, then the two cycles are of infinite order for any $e$ [5].
@@
    Infinitude of order in the Chow group: Using the work of Darmon-Rotger-Sols [8] on Chow-Heegner points, the first-named author and K. Murty showed in [9] that for every positive integer $N$ that has a prime divisor greater than 7, the Ceresa cycle of the Fermat curve $F(N)$ of degree $N$ is non-torsion in the Chow group for every choice of $e$. This work was generalized by the second author in [20] to an infinite number of non-hyperelliptic simple Fermat quotients. [^2] In [9] and [20] $e$ was a point of the curve, but the proofs also work if $e$ is a divisor of degree 1. Using the same method, Kerr-Li-Qiu-Yang [14] and Lupoian-Rawson [19] have proved that the Ceresa cycles of an infinite number of modular curves are non-torsion in the Chow group.
@@
We prove two main results. The first result is the following (see Theorem 4.2 for a slightly more general statement):
@@
Let $p\geq 3$ be a prime and $e\in X_{p,\lambda}$ be a cusp (i.e., a point satisfying $x_{1}x_{2}y_{1}y_{2}=0$). Then for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$ and positive integer $l\leq(p-1)/2$, we have
@@
where $\Phi_{1}(\Delta_{{\rm GKS},e}(X_{p,\lambda}))$ is the complex Abel-Jacobi image of $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ in the intermediate Jacobian of the integral Hodge structure $H_{3}(X_{p,\lambda}^{3})$. In particular, the modified diagonal cycle $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ is nontrivial in the integral Chow group $\operatorname{CH}_{1}(X_{p,\lambda}^{3})$ for every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$.
It has been observed in [16, Proposition 2.2] that given any smooth projective complex curve $X$ of genus $g$ with canonical divisor $K(X)$, if $e$ is a degree 1 divisor on $X$ such that $(2g-2)[e]\neq K(X)$ in $\operatorname{CH}_{0}(X)\otimes\mathbb{Q}$, then the modified diagonal cycle of $X$ at $e$ is non-torsion in the Chow group. For a cuspidal choice of a base point $e$ as in Theorem 1.1, one has $(2(p-1)^{2}-2)[e]=K(X_{p,\lambda})$ in $\operatorname{CH}_{0}(X_{p,\lambda})\otimes\mathbb{Q}$ (see Corollary 3.4). Thus Theorem 1.1 does not follow from [16, Proposition 2.2].
To our knowledge, Theorem 1.1 provides the first explicit example of a smooth non-constant family $\mathscr{X}\rightarrow S$ of curves of arbitrarily large genus over a base $S$ where the modified diagonal cycle is nontrivial in the entire family. Since the hypergeometric curve $X_{p,\lambda}$ is defined over $\mathbb{Q}(\lambda)$, in particular, the result gives infinitely many curves of a given genus $(p-1)^{2}$ (for every prime $p\geq 3$) over any subfield of $\mathbb{C}$ with nontrivial modified diagonal cycles. By the recent result [18, Proposition 5.6], Theorem 1.1 also implies the nontriviality of the Ceresa cycles of hypergeometric curves $X_{p,\lambda}$ for $p\geq 5$ at cuspidal base points (see Example 5.15 in loc. cit.).
Unfortunately, we do not know whether $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ is non-torsion or not for an arbitrary $p$. However, for $p=3$ we can prove the following:
@@
For every $\lambda\in\mathbb{P}^{1}\setminus\{0,1,\infty\}$ and every choice of a base divisor $e$ of degree 1 on $X_{3,\lambda}$ supported on the cusps, the modified diagonal cycle $\Delta_{{\rm GKS},e}(X_{3,\lambda})$ (and hence the Ceresa cycle $\operatorname{Cer}_{e}(X_{3,\lambda})$) is torsion in the Chow group.
The curve $X_{3,\lambda}$ is not hyperelliptic (see Proposition 3.1), so this gives an explicit example of a family of non-hyperelliptic curves of genus 4 with torsion modified diagonal cycles. Another example of such a family has been given by Qiu-Zhang [25]. The proof of Theorem 1.2 uses the cohomological vanishing criteria of Qiu-Zhang [25, Theorem 1.2.1] and Laga-Shnidman [17, Theorem A]. [^4] We use the latter criterion to make the argument shorter, but the former criterion would also suffice in this case. We expect that for $p\geq 5$ and a cuspidal divisor $e$, the cycle $\Delta_{{\rm GKS},e}(X_{p,\lambda})$ is of infinite order for “most” (or perhaps all) $\lambda$.
@@ 2.2. A criterion for finiteness of the order of the modified diagonal cycle
Let $X$ be a smooth projective curve of genus $g\geq 3$ over $\mathbb{C}$. Let $\operatorname{Jac}(X)$ and $K(X)$ be respectively the Jacobian and canonical divisor of $X$. Given a divisor $e$ of degree 1 on $X$, let
@@
be the modified diagonal cycle defined in [11] (see the §1 to recall the definition). We refer to this as the Gross-Kudla-Schoen modified diagonal cycle with base divisor (or point, if $e$ is a point) $e$. Gross and Schoen showed that this cycle is homologically trivial [11, Proposition 3.1]. One also has the Ceresa cycle
@@
(see §1). By [16, Proposition 2.2], if $\Delta_{{\rm GKS},e}(X)$ (resp. $\operatorname{Cer}_{e}(X)$) is torsion in the Chow group, then (2g2)eK(X)(2g-2)e-K(X) is a torsion point of $\operatorname{Jac}(X)\cong\operatorname{CH}_{0}(X)_{\mathrm{hom}}$. Thus the question of whether $\Delta_{{\rm GKS},e}(X)$ (resp. $\operatorname{Cer}_{e}(X)$) is torsion in the Chow group is reduced to the case when (2g2)e=K(X)(2g-2)e=K(X) in $\operatorname{CH}_{0}(X)_{\mathrm{hom}}\otimes\mathbb{Q}$. For such a base divisor $e$, by [27, Theorem 1.5.5], the finiteness of the orders of $\operatorname{Cer}_{e}(X)$ and $\Delta_{{\rm GKS},e}(X)$ in the corresponding Chow groups are equivalent. Qui and W. Zhang [25, Theorem 1.2.1] have given a sufficient cohomological criterion for this finiteness. This criterion has been strengthened by Laga and Shnidman in [17, Theorem A]. We recall the latter criterion, stated for $\Delta_{{\rm GKS},e}(X)$ in terms of the de Rham cohomology:
Theorem 2.1 (Laga and Shnidman [17]).
Let $X$, $g$, and $K(X)$ be as above. Let $e$ be a divisor of degree 1 on $X$ such that (2g2)e=K(X)(2g-2)e=K(X) in $\operatorname{CH}_{0}(X)_{\mathrm{hom}}\otimes\mathbb{Q}$. Let $\operatorname{Aut}(X)$ be the automorphism group of $X$. Consider the action of $\operatorname{Aut}(X)$ on $\wedge^{3}H^{1}_{\rm dR}(X)$ through its natural action on each factor of the de Rham cohomology $H^{1}_{\rm dR}(X)$, and let $(\wedge^{3}H^{1}_{\rm dR}(X))^{\operatorname{Aut}(X)}$ be the fixed part under this action. If $(\wedge^{3}H^{1}_{\rm dR}(X))^{\operatorname{Aut}(X)}=0$, then $\Delta_{{\rm GKS},e}(X)$ is torsion in the Chow group.
Note that [17] works with Chow groups with coefficients in $\mathbb{Q}$, hence the assertion is stated as a vanishing statement in that paper. Also note that [17, Theorem A] is slightly stronger than the statement recorded above (and needed in this paper), as Laga and Shnidman only require the vanishing of the subspace of the primitive part of $\wedge^{3}H^{1}_{\rm dR}(X)=H^{3}_{\mathrm{dR}}(\operatorname{Jac}(X))$ fixed by $\operatorname{Aut}(X)$.
@@ 2.3. Hypergeometric curves
The differential form $\omega^{a,b}_{N,\lambda}$ is holomorphic on $X_{N,\lambda}$, and $\eta^{a,b}_{N,\lambda}$ is non-holomorphic but of the second kind (= meromorphic with zero residues everywhere). The collection of differential forms $\omega^{a,b}_{N,\lambda}$ and $\eta^{a,b}_{N,\lambda}$ forms a basis of $H^{1}_{\rm dR}(X_{N,\lambda})^{\chi_{N}^{a,b}}$.
@@ 3.1. Determination of the hyperelliptic cases
In this subsection we will determine when a hypergeometric curve is hyperelliptic. This is important to put Theorems 1.1 and Theorem 1.2 in a better context, since the behaviour of the modified diagonal cycle for hyperelliptic curves is well understood. We also determine when a quotient $C_{N,\lambda}^{a,b}$ of a hypergeometric curve is hyperelliptic. This latter result will not be used in the proofs of Theorems 1.1 and 1.2, however, it may be helpful for future study of modified diagonal cycles of these quotients. The result is also relevant to Proposition 4.4 in §4.3.
@@
Suppose that $X_{N,\lambda}$ is hyperelliptic. Let $\tau$ be its hyperelliptic involution. Since the genus of $X_{N,\lambda}$ is greater than one, $\tau$ lies in the center of the automorphism group of $X_{N,\lambda}$, so that we have
$\operatorname{Aut}(X_{N,\lambda})/\langle\tau\rangle\subset\operatorname{Aut}(X_{N,\lambda}/\tau)=\operatorname{Aut}(\mathbb{P}^{1})=\operatorname{PGL}_{2}(\mathbb{C}).$
First assume that $\tau\not\in\mu_{N}\times\mu_{N}$. Then we have $\mu_{N}\times\mu_{N}\subset\operatorname{Aut}(X_{N,\lambda})/\langle\tau\rangle$. By the well known characterization of the finite subgroups of $\operatorname{PGL}_{2}(\mathbb{C})$ due to Klein, $\operatorname{Aut}(X_{N,\lambda})/\langle\tau\rangle$ is isomorphic to one of the following: the cyclic group $C_{n}$ of order $n\geq 1$, the dihedral group $D_{n}$ of order $2n$ for $n\geq 2$, the alternating group $A_{4}$ or $A_{5}$, or the symmetric group $S_{4}$. However, these groups do not include any copy of the group $\mu_{N}\times\mu_{N}$.
Next, we assume that $\tau\in\mu_{N}\times\mu_{N}$. Then $N$ is even and $\tau$ is given by one of the three elements of order 2 in $\mu_{N}\times\mu_{N}\subset\operatorname{Aut}(X_{N,\lambda})$, i.e., one of the following:
@@
We now consider the quotients of $X_{N,\lambda}$.
@@
First, we assume that $a=b$ or $a+b=N$. By (2.3) and (2.4), it suffices to show the case $(a,b)=(1,1)$ and (1,N1)1,N-1). There are isomorphisms